OpenStudy (anonymous):

You pay off a $10,000 loan with an APR of 12% over 10 years with fixed monthly payments a)how big are the payments b) how much do you pay total c)what fraction of what you paid goes to interest?

OpenStudy (dumbcow):

use finance calculator like a BA II A) enter 120 for n (num of months) enter 1% for monthly int rate enter 10,000 for present value press compute payment payment = 143.47 B) total = payment*num months total = 143.47*120 = 17,216.40 C) (total - loanAmount) /loanAmount = (17,216.40-10,000)/10,000 = 72.16%